# Skill-informed Data-driven Haptic Nudges for High-dimensional Human Motor Learning

### 13 Mar 2026 v1

@@ 5 Results
This section presents the findings from our human-subject study, designed to test the hypothesis that haptic feedback informed by the learner’s latent skill state accelerates motor acquisition and enhances task performance relative to both a no-feedback baseline and a standard heuristics-based strategy.
@@ 5.1 Experiment Details
To validate our approach, we divided participants into three distinct experimental groups, all of whom performed the target capture task described in Section 2.1. These groups were defined as follows: a Control Group, which received no haptic feedback; a Heuristics Group, which received feedback based on a heuristics-based nudge feedback policy; and a POMDP Group, which received feedback derived from our optimal, skill-state-aware policy. Each group consisted of $10$ participants [^3] All human subject experiments were approved under the Michigan State University Institutional Review Board Study ID LEGACY14-431M.. The experimental timeline was structured sequentially. First, data was collected from the Heuristics Group. The input-output gameplay data from the initial five participants in this group were utilized to train the Input-Output Hidden Markov Model (IOHMM), establishing the baseline behavioral model of human motor learning. For the POMDP Group, the haptic feedback was generated by casting the problem as a POMDP and solving it approximately via the QMDP method. During the experiment, the system maintained a running belief state, which was updated after every trial. To select an optimal action for the subsequent trial, a state was sampled from this current belief distribution and mapped to a specific nudge finger index via the QMDP policy. The belief state was initialized using the prior distribution $\mathbb{P}(h_{0})$ derived from the trained IOHMM.
@@
5.2 Task Performance: Straightness of Trajectory
We used the Straightness of Trajectory SoT to measure task performance across the three groups, where a lower value signifies a smoother, more desirable movement. We conducted two primary analyses to compare the three participant groups. First, we tracked the evolution of SoT across all game trials to assess the overall performance trajectories. Second, we computed the number of trials required for participants to converge to a predefined SoT threshold. This convergence analysis directly evaluates the efficiency of a nudging policy, as haptic feedback can nudge participants towards efficient joint motions that could improve task performance and thus require fewer trials to reach a performance threshold. Moreover, a nudging policy based on the motor skill state of the participant could help achieve accelerated motor skill acquisition by optimally nudging the right joints towards better synergistic patterns, going beyond simple task performance improvements. To analyze performance across the full learning trajectory, we fit a Linear Mixed Model (LMM) to the SoT values for all $480$ gameplay trials from each participant, with trial number, group, and their interaction as fixed effects, and a random intercept per participant, estimated via Restricted Maximum Likelihood (REML) [corbeil1976restricted]. This model accounts for the correlation structure arising from repeated measurements within participants. Post-hoc two-tailed contrasts with Bonferroni correction were used to test for pairwise differences between the POMDP, heuristic, and control groups. For the threshold-based analysis, where each participant contributes a single scalar value (the trial index at which a given SoT threshold is first reached), we conducted a one-way ANOVA to test for an overall group effect, followed by pairwise independent-samples t-tests (two-tailed) with Bonferroni correction.
The results from Fig. 4 show that SoT decreases with trials for all groups, and the mean number of trials taken to achieve a particular SoT threshold also decreases for increasing thresholds, indicating task learning improvement through training. The p-values from Table 1 show that participants trained on the optimal nudging policy differ statistically significantly (at a significance level of $\alpha=0.10$) in average SoT convergence across the game trials as compared to participants trained without any haptic feedback. On average, the POMDP group takes fewer trials than the control group to converge to different SoT thresholds. This indicates expedited performance improvement in the early stages of training. Fig. 4 shows that the POMDP group takes fewer trials to hit SoT thresholds compared to the heuristic group, and has faster SoT convergence throughout the game trials; however, the results are not statistically significant.
Figure 5: Motor Output Performance: (a) Mean RE curves (with $95\%$ confidence intervals) of the three group participants, and (b) the mean number of trials (along with the scatter plot taken across participants) to converge to different RE thresholds under the three nudging policies shows fastest convergence for participants trained on POMDP-based optimal nudging policy. The low two-tailed test p-values from the linear mixed model fits show a decrease in RE across trials and in the number of trials required by the POMDP group for all RE thresholds. Each scatter plot point represents the number of trials taken by a participant to achieve that RE threshold.
@@ 5.2 Task Performance: Straightness of Trajectory
5.3 Motor Output Performance: Reaching Error
We used the reaching error RE to measure the motor output performance for the three experiment groups. Recall that the RE for a game trial is the Euclidean distance of the screen cursor from the trial target at the end of 22~s, or end of movement, whichever is earlier. Thus, lower RE values signify better motor output and task learning. We performed two analyses: first, we studied the evolution of RE across gameplay trials to assess overall motor performance, and second, we tracked the number of trials required for participants to converge to predefined RE thresholds. The statistical analyses employed here are the same as in Section  5.2.
Results from Fig. 5 show that the RE decreases across all groups as the gameplay progresses, indicating improved motor performance with training. The POMDP group achieves statistically significantly (at significance level $\alpha<0.10$) lower RE compared to the other two groups across all gameplay trials, as shown in Table 2. Results also show that the POMDP group requires fewer trials on average to converge to lower RE thresholds (Fig. 5(b)) than the control and the heuristic group, indicating expedited motor performance improvement; however, the results are not statistically significant.
@@ 5.3 Motor Output Performance: Reaching Error
5.4 Synergy Analysis
The inherent high redundancy in the target capture game requires participants to idenepsy an underlying manifold to effectively map finger joint movements to desired cursor motion towards the targets. To characterize how coordination strategies emerge, we analyzed the simplification of motor-control dimensionality across the three groups by calculating the Variance Accounted For (VAF) by the first few Principal Components (PCs) derived from the hand postures data (finger joint angles $\boldsymbol{q}$). For this analysis, the eight training blocks were condensed into four phases (two blocks per phase) for ease of representation. We also analyze the number of PCs required to explain more than $90\%$ variance in the data across the three groups. We hypothesized that the POMDP group would exhibit a faster increase in top-PC VAF, indicating more efficient discovery of the 22D task-relevant subspace.
Group Comparison: As illustrated in Fig. 6, the POMDP group achieved higher average VAF in the first four PCs, particularly in later training phases, and required fewer PCs on average to explain $90\%$ of the variance in joint angle data. These results provide a structural explanation for the POMDP group’s improved task learning and motor output performance. By discovering the task space more rapidly than other groups, POMDP group participants successfully coordinated their $20$ finger joints into a simpler, more effective strategy, which is a direct indication of accelerated skill acquisition driven by the optimal nudge feedback.
@@ 5.5 Learning Dynamics and Convergence Analysis
To quanepsy the temporal dynamics of motor learning beyond simple performance metrics, we analyzed the evolution of the participants’ latent states throughout the experiment. While the IOHMM provides a discrete set of hidden states representing different performance strategies, the POMDP framework maintains a continuous belief state $b_{k}\in\mathbb{R}^{7}$ representing the probability distribution over these states at trial $k$. This probabilistic approach allows us to capture the uncertainty and the transitional phases of learning that discrete classification might miss.
Expected Latent State: The state ordering obtained in Section 3.3 is preserved in this analysis, with state 0 as the best and state $6$ as the worst performing. To visualize the learning trajectory, we computed the Expected Latent State for each trial, defined as the probability-weighted average of the state ranks as
$\mathbb{E}[s_{k}]=\sum_{i=0}^{6}i\cdot b_{k}(i)$ (13)
where $b_{k}(i)$ is the belief (probability) that the participant is in the $i$-th ranked state at trial $k$. A value of $\mathbb{E}[s_{k}]$ closer to 0 indicates a high probability of having adopted an optimal control strategy.
Convergence Results: Figure 7 illustrates the average Expected State trajectories for the three experimental groups over the course of the $480$ trials. Shaded regions indicate the standard error of the mean. All groups exhibited a characteristic learning curve, initiating the experiment with an Expected State around $5$, indicating initial exploratory behavior and high motor error. As the experiment progressed, a monotonic mean decrease in $\mathbb{E}[s_{k}]$ was observed, reflecting the acquisition of the BoMI mapping $C$.
Group Comparison: The POMDP group demonstrated an expedited rate of convergence compared to the other two groups. Specifically, the POMDP group reached a steady-state mastery level (defined as $\mathbb{E}[s_{k}]<3$) by trial $150$ (median), whereas the control and the heuristic groups reached this threshold statistically significantly later ($p<0.0001$ for Trials(POMDP) << Trials(Control), and $p=0.0064$ for Trials(POMDP) << Trials(Heuristic)). This suggests that the optimal nudging policy effectively guided participants away from suboptimal local minima and facilitated a more rapid transition into high-performance states. The lower entropy in the belief distribution for the POMDP group towards the end of the session further indicates that these participants not only performed better but had settled into a more stable and confident control strategy compared to other groups. Since the trials to reach this threshold were not normally distributed across the group participants, we used the Kruskal-Wallis test for statistical significance (at significance level $\alpha<0.05$). This was followed by post-hoc one-tailed contrasts to compare the POMDP group with the other two groups.
@@ 6 Discussion
This work proposes a POMDP-based framework for expediting complex motor skill acquisition and task performance using skill-state-based optimal haptic feedback policy. This proposed approach shows accelerated task performance compared to no feedback, and a performance heuristics-based feedback design by leveraging the learner’s estimated motor skill state. These promising results underscore the significant potential of skill-informed haptic feedback in improving the motor learning process and the corresponding task performance, particularly in novel complex motor skill acquisition tasks.
@@ 6.1 Decoupling Latent Skill from Observable Performance
A central pillar of our approach is the distinction between task/motor performance and latent skill acquisition. Our results demonstrate that nudge feedback policies targeting observable performance metrics (Heuristic Group) do not necessarily lead to superior learning and performance outcomes compared to no-feedback baselines (Control Group). This aligns with the ’Guidance Hypothesis’ in motor learning literature [7], which suggests that outcome-based error-correction can engender dependency, allowing the learner to perform well during training without internalizing the task dynamics.
In contrast, the POMDP-based optimal nudging policy effectively decoupled execution noise from the underlying true skill through the estimated behavioral model of motor learning. By optimizing for long-term state transitions ($C_{\textup{gen}}$) rather than immediate motor performance ($C_{\textup{imm}}$), the optimal policy likely permitted short-term performance deviations if they facilitated exploration of the state space. This supports the notion that effective robotic tutoring should operate as an active teacher, prioritizing the resolution of internal model inaccuracies/uncertainties over instantaneous kinematic error corrections. The statistical superiority of the POMDP group in both RE and SoT performance and convergence (Tables 1 and 2) validates the utility of modeling learning as evolution through discrete latent states rather than a continuous improvement in task performance.
@@ 6.2 Accelerated Discovery of Better Task Manifolds
High-dimensional de-novo tasks require learners to solve the redundancy problem: idenepsying a low-dimensional subspace (manifold) within the high-dimensional task space that maps effectively to the task goal. Our synergy analysis (Fig. 6) reveals that the POMDP group exhibited a more rapid consolidation of joint space variance into the first few PCs compared to the other two groups.
This suggests that the skill-state aware nudge feedback did not merely improve task performance, but actively guided participants toward the discovery of the better task manifolds. This behavior is also supported by the learning dynamics and convergence analysis results in Fig. 7. Critically, by penalizing the transition to low-performing skill states, the POMDP policy likely encouraged the utilization of correct fingers sequentially to foster coordinated joint movement strategies toward expedited skill acquisition. This finding is consistent with the Uncontrolled Manifold (UCM) hypothesis, suggesting that optimal assistance facilitates the separation of task-relevant and task-irrelevant variabilities, facilitating the discovery and utilization of effective synergies.
@@ 6.3 Interpretability of Data-driven Behavior Models
This aligns with the kinematic realities of our setup and the BoMI, where the thumb possesses a larger range of motion and greater contribution to cursor motions than the other fingers. The model learned these biomechanical constraints solely from the input-output data, without explicit biomechanical modeling. This validates the IOHMM not just as a control tool, but as a diagnostic tool capable of idenepsying which fingers are most receptive to haptic interventions at specific stages of learning. Such interpretability is crucial for clinical application, where therapists must trust the agent to prescribe physiologically valid movement patterns.
@@ 6.4 Limitations and Future Directions
While our results are promising, several limitations warrant discussion. First, the discretization of skill state into $N=7$ states is an abstraction of the continuous motor learning process. While necessary for the tractability of the QMDP solver, this quantization may mask subtle changes/adjustments in a strategy that occur within a single state. Future work could explore continuous state POMDP solvers or hybrid approaches to capture the finer details/granularities in the learning dynamics.
Second, our study validated the proposed framework on healthy humans learning a novel motor task. While this paradigm is an established proxy for the motor re-learning process, post neurological injury populations introduce state-dependent constraints owing to specific impairments (eg, spasticity, paresis) that are not modeled in healthy participants. However, the data-driven nature of our IOHMM framework is agnostic to specific learning dynamics. In principle, the same pipeline could be applied to patient data to learn the personalized transition model; however, this would necessitate the use of individualized models instead of a population-average model. Another possibility is to estimate a population average model, personalizing it to each individual using real-time input-output data during the gameplay. Future efforts could focus on transitioning this framework to clinical cohorts to validate its efficacy in neuro-rehabilitation.

### 13 Apr 2026 v2

@@ 5 Results
This section presents the findings from our human-subject study, designed to evaluate the efficacy of our proposed optimal nudging policy compared to a no-feedback baseline and a standard heuristics-based strategy.
@@ 5.1 Experiment Details
To validate our approach, we divided participants into three distinct experimental groups, all of whom performed the target capture task described in Section 2.1. These groups were defined as follows: a Control Group, which received no haptic feedback; a Heuristic Group, which received feedback based on the heuristic nudge feedback policy; and a POMDP Group, which received our optimal, skill-state-aware nudge feedback. Each group consisted of $10$ participants [^2] All human subject experiments were approved under the Michigan State University Institutional Review Board Study ID LEGACY14-431M.. The experimental timeline was structured sequentially. First, data was collected from the Heuristic Group. The input-output experiment data from the initial five participants in this group were utilized to train the Input-Output Hidden Markov Model (IOHMM), establishing the baseline behavioral model of human motor learning. For the POMDP Group, the haptic feedback was generated by casting the problem as a POMDP and solving it approximately via the QMDP method. During the experiment, the system maintained a running belief state, which was updated after every trial. To select an optimal action for the subsequent trial, a state was sampled from this current belief distribution and mapped to a specific nudge finger index via the QMDP policy. The belief state was initialized using the prior distribution $\mathbb{P}(h_{0})$ derived from the trained IOHMM.
5.2 Movement Efficiency: Straightness of Trajectory
We used the Straightness of Trajectory SoT to measure movement efficiency across the three groups, where a lower value signifies a smoother, more desirable movement. We conducted two primary analyses to compare the three participant groups. First, we tracked the evolution of SoT across all experiment trials to assess the overall performance trajectories. Second, we computed the number of trials required for participants to converge to a predefined SoT threshold. This convergence analysis directly evaluates the efficiency of a nudging policy, as haptic feedback can nudge participants towards efficient joint movements that could improve task performance and thus require fewer trials to reach a performance threshold. Moreover, a nudging policy based on the motor skill state of the participant could help achieve accelerated motor skill acquisition by optimally nudging the right joints towards better synergistic patterns, going beyond simple task performance improvements. To analyze performance across the full learning trajectory, we fit a Linear Mixed Model (LMM) to the SoT values for all $480$ experiment trials from each participant, with trial number, group, and their interaction as fixed effects, and a random intercept per participant, estimated via Restricted Maximum Likelihood (REML) [corbeil1976restricted]. This model accounts for the correlation structure arising from repeated measurements within participants. Post-hoc two-tailed contrasts with Bonferroni correction were used to test for pairwise differences between the POMDP, heuristic, and control groups. For the threshold-based analysis, where each participant contributes a single scalar value (the trial index at which a given SoT threshold is first reached), we conducted a one-way ANOVA to test for an overall group effect, followed by pairwise independent-samples t-tests (two-tailed) with Bonferroni correction.
The results from Fig. 4 show that SoT decreases with trials for all groups, and the mean number of trials taken to achieve a particular SoT threshold also decreases for increasing thresholds, indicating task learning improvement through training. The p-values and Cohen’s d-values from Table 1 show that participants trained on the optimal nudging policy achieved lower SoT values on average across the experiment trials as compared to participants trained with no haptic nudges and heuristic nudges; however, the results are not statistically significant. Furthermore, the POMDP group takes statistically significantly (at significance level $\alpha=0.10$) fewer trials than the control group to converge to higher SoT thresholds. This indicates expedited performance improvement in the early stages of training. Although the POMDP group converged to various SoT thresholds earlier than the heuristic group on average, the results are not statistically significant.
@@ 5.2 Movement Efficiency: Straightness of Trajectory
Figure 5: Endpoint Accuracy: (a) Mean RE curves (with $95\%$ confidence intervals) of the three group participants, and (b) the mean number of trials (along with the scatter plot taken across participants) to converge to different RE thresholds under the three nudging policies shows fastest convergence for participants trained on POMDP-based optimal nudging policy. The low two-tailed test p-values from the linear mixed model fits show a decrease in RE across trials and in the number of trials required by the POMDP group for all RE thresholds. Each scatter plot point represents the number of trials taken by a participant to achieve that RE threshold.
@@
5.3 Endpoint Accuracy: Reaching Error
We used the reaching error RE to measure the endpoint accuracy of the three experiment groups. Recall that the RE for an experiment trial is the Euclidean distance of the screen cursor from the trial target at the end of 22~s, or end of movement, whichever is earlier. Thus, lower RE values signify better task performance. We performed two analyses: first, we studied the evolution of RE across experiment trials to assess overall motor performance, and second, we tracked the number of trials required for participants to converge to predefined RE thresholds.
Results from Fig. 5 show that the RE decreases across all groups as the experiment progresses, indicating improved motor performance with training. The POMDP group achieves statistically significantly (at significance level $\alpha<0.10$) lower RE compared to the other two groups across all experiment trials, as shown in Table 2. Results also show that the POMDP group requires fewer trials on average to converge to lower RE thresholds (Fig. 5(b)) than the control and the heuristic group, indicating expedited motor performance improvement; however, the results are not statistically significant.
5.4 Synergy Analysis
The inherent high redundancy in the target capture task requires participants to identify an underlying manifold to effectively map finger joint movements to desired cursor motion towards the targets. To characterize how coordination strategies emerge, we analyzed the simplification of motor-control dimensionality across the three groups by calculating the Variance Accounted For (VAF) [17] by the first few Principal Components (PCs) derived from the hand postures data (finger joint angles $\boldsymbol{q}$). For this analysis, the eight training blocks were condensed into four phases (two blocks per phase) for ease of representation. We also analyze the number of PCs required to explain more than $90\%$ variance in the data across the three groups. We hypothesized that the POMDP group would exhibit a faster increase in top-PC VAF, indicating more efficient discovery of the 22D task-relevant subspace.
Group Comparison: As illustrated in Fig. 6, the POMDP group achieved higher average VAF in the first four PCs, particularly in later training phases, and required fewer PCs on average to explain $90\%$ of the variance in joint angle data, although the results are not statistically significant. These results provide a structural explanation for the POMDP group’s improved task learning and motor performance. By discovering the task space more rapidly than other groups, POMDP group participants successfully coordinated their $20$ finger joints into a simpler, more effective strategy, which is a direct indication of accelerated skill acquisition driven by the optimal nudge feedback.
@@ 5.5 Learning Dynamics and Convergence Analysis
To quantify the temporal dynamics of motor learning beyond simple performance metrics, we analyzed the evolution of the participants’ latent states throughout the experiment. The POMDP framework maintains a continuous belief state $b_{k}\in\Delta_{7}$ representing the probability distribution over the IOHMM states at trial $k$.
Expected Latent State: To visualize the learning trajectory, we computed the Expected Latent State for each trial, defined as the probability-weighted average of the state ranks as
$\mathbb{E}[s_{k}]=\sum_{i=0}^{6}i\cdot b_{k}(i),$ (15)
where $b_{k}(i)$ is the belief (probability) that the participant is in the $i$-th ranked state at trial $k$. The smaller values of $\mathbb{E}[s_{k}]$ indicate a high probability of having adopted an optimal control strategy.
Convergence Results: Fig. 7 illustrates the distribution of the average expected state for the three experimental groups, aggregated across four phases (two training blocks per phase). All groups exhibited a characteristic learning progression, initiating Phase $1$ with high expected states, indicating initial exploratory behavior and high motor error. As the experiment progressed, a monotonic decrease in median $\mathbb{E}[s_{k}]$ was observed across all groups, reflecting the acquisition of the BoMI mapping $C$.
Group Comparison: The POMDP group demonstrated a markedly expedited rate of convergence compared to the other two groups. Notably, during Phase $1$, the POMDP group already achieved a median expected state below a pre-defined mastery threshold (defined as $\mathbb{E}[s_{k}]<3$), whereas the other two groups maintained median states above this threshold. To quantify this temporal advantage, we evaluated the specific number of trials required for participants to reach this steady-state mastery level. The POMDP group reached mastery by trial $150$ (median), whereas the other two groups reached this threshold statistically significantly later ($p<0.0001$ for Trials(POMDP) << Trials(Control), and $p=0.0064$ for Trials(POMDP) << Trials(Heuristic)). This suggests that the optimal nudging policy effectively guided participants away from suboptimal local minima and facilitated a more rapid transition into high-performance states. The lower entropy in the belief distribution for the POMDP group towards the end of the session further indicates that these participants not only performed better but had settled into a more stable and confident control strategy compared to other groups. Since the trials to reach this threshold were not normally distributed across the group participants, we used the Kruskal-Wallis test for statistical significance (at significance level $\alpha<0.05$). This was followed by post-hoc one-tailed contrasts to compare the POMDP group with the other two groups.
@@ 6 Discussion
This work proposes a POMDP-based framework for expediting complex motor skill acquisition and task performance using skill-state-based optimal haptic feedback policy. This proposed approach shows accelerated task performance compared to no feedback and a heuristic feedback design by leveraging the learner’s estimated motor skill state. These promising results underscore the significant potential of skill-informed haptic feedback in improving the motor learning process and the corresponding task performance, particularly in novel complex motor skill acquisition tasks.
@@ 6.1 Decoupling Latent Skill from Observable Performance
A central pillar of our approach is the distinction between task performance and latent skill acquisition. Our results demonstrate that nudge feedback policies targeting simple heuristic policies do not necessarily lead to superior learning and performance outcomes compared to no-feedback baselines (Control Group). This aligns with the guidance hypothesis in motor learning literature [13], which suggests that outcome-based error-correction can engender dependency, allowing the learner to perform well during training without internalizing the task dynamics.
In contrast, the POMDP-based optimal nudging policy unified multiple performance measures into a unified skill state while accounting for the motor noise. By optimizing for long-term cumulative reward, with reward for each trial designed to promote better performance on the current trial as well as all possible task types in the next trial,, the optimal policy likely permitted short-term performance deviations if they facilitated effective exploration of motor space. This supports the notion that effective robotic tutoring should operate as an active teacher, prioritizing the improvement in latent skill state over instantaneous endpoint error corrections.
@@ 6.2 Accelerated Discovery of Better Task Manifolds
High-dimensional de-novo tasks require learners to solve the redundancy problem: identifying a low-dimensional subspace within the high-dimensional task space that maps effectively to the task goal. Our synergy analysis (Fig. 6) reveals that the POMDP group exhibited a more rapid consolidation of joint space variance into the first few PCs compared to the other two groups.
This suggests that the skill-state aware nudge feedback did not merely improve task performance, but actively guided participants toward the discovery of the better joint movement subspaces. This behavior is also supported by the learning dynamics and convergence analysis results in Fig. 7. Critically, by penalizing the transition to low-performing skill states, the POMDP policy likely encouraged the utilization of correct fingers sequentially to foster coordinated joint movement strategies toward expedited skill acquisition. This finding is consistent with the Uncontrolled Manifold (UCM) hypothesis, suggesting that optimal assistance facilitates the separation of task-relevant and task-irrelevant variabilities, facilitating the discovery and utilization of effective synergies.
@@ 6.3 Interpretability of Data-driven Behavior Models
This aligns with the kinematic realities of our setup and the BoMI, where the thumb possesses a larger range of motion and greater contribution to cursor motions than the other fingers. The model learned these biomechanical constraints solely from the input-output data, without explicit biomechanical modeling. This validates the IOHMM not just as a control tool, but as a diagnostic tool capable of identifying which fingers are most receptive to haptic interventions at specific stages of learning. Such interpretability is crucial for clinical application, where therapists must trust the agent to prescribe physiologically valid movement patterns.
@@ 6.4 Limitations and Future Directions
While our results are promising, several limitations warrant discussion. First, the discretization of skill state into $N=7$ states is an abstraction of the continuous motor learning process. While necessary for the tractability of the QMDP solver, this quantization may mask subtle changes/adjustments in a strategy that occur within a single state. Future work could explore continuous state POMDP solvers or hybrid approaches to capture the finer details/granularities in the learning dynamics. However, data constraints may make fitting a continuous state model reliably a challenging task.
Second, our study validated the proposed framework on healthy humans learning a novel motor task. While this paradigm is an established proxy for the motor re-learning process, post-neurological injury populations introduce state-dependent constraints owing to specific impairments (eg, spasticity, paresis) that are not modeled in healthy participants. However, the data-driven nature of our IOHMM framework is agnostic to specific learning dynamics. In principle, the same approach could be applied to patient data to learn the personalized transition model; however, this would necessitate the use of individualized models instead of a population-average model. Another possibility is to estimate a population average model, personalizing it to each individual using real-time input-output data during the experiment. Future efforts could focus on transitioning this framework to clinical cohorts to validate its efficacy in neuro-rehabilitation.